Two investments totaling $27,500 produce an annuel income of $1,895. One investment yields 7% a year, while the other yields 4% per year. How much is invested at each rate?
Answer by mananth(16946)

Investment Part I 7.00% per annum ----x
Investment part II 4.00% per annum ----y
The sum of the investments is $27,500.00
The sum of individual interests = $1,895.00
x+y=27500 ------------------------1
7.00%x+4.00%y=$1,895.00
Multiply by 100 ( to eliminate the %)
7x+4y=$189,500.00 --------2

Multiply (1) by -7
we get

-7x-7y =-192500.00

Add this to (2)

-3y=-$3,000.00

divide by-3

y = $1,000.00 investment at 4.00%
Balance $26,500.00 investment at 7.00%
CHECK
$26,500.00 @ 7.00% $1,855.00
$1,000.00 @ 4.00% $40.00
Total -------------------- $1,895.00